# SPARTAN: A Sparse Transformer Learning Local Causation

### v2

@@ 3.2 Interventions
We can extend our model to represent interventions in the cases where the model is trained on a dataset with a set of environments with intervened dynamics. During training, the model has access to the environment index $I$ for each transition but does not know which object is affected by the intervention. Conditioned on the current state $s^{t}_{1:N}$ and the environment index $I$, the model needs to identify a subset of $s^{t+1}_{1:N}$ as intervention targets and change the predictions (cf. Eq. 2). To do this, we keep a learnable codebook of intervention tokens, $\mathcal{T}_{1:K}\in\mathbb{R}^{d\times k}$, where $d$ is the dimension of the state tokens and $K$ is the number of interventions. We append the corresponding intervention token $\mathcal{T}_{I}$ to the object tokens and proceed as described in the previous subsection. Here, $\bar{A}_{i(N+1)}=0$ indicates that there is no path from the intervention token to $s^{t+1}_{i}$ and therefore $s^{t+1}_{i}$ is not an intervention target, i.e. the prediction for $s^{t+1}_{i}$ is not affected to the changes in this environment.
@@ 3.3 Training
As discussed in Sec. 2.2, our aim is to fit the data distribution with the sparsest possible model in terms of causal edges and intervention targets. The training objective is to minimise the expected model loss with sparsity regularisation,
@@
where $\theta$ are the model parameters, including the parameters for the Transformer and the intervention tokens, $\mathcal{L}$ is a loss function such as MSE, and $\lambda_{s}$ denotes the regularisation hyperparameter. Note that due to the intervention token being part of the input during training, $|\bar{A}|$ is the sum of the number of causal edges and the intervention targets. Hence this objective is analogous to Eq. 3 except we penalise the expected $|\bar{A}|$.
In practice, the model is sensitive to the choice of $\lambda$ since a high $\lambda$ leads to mode collapse where the model has no connections. Since removing non-causal edges should not deteriorate the prediction accuracy, we formulate a constrained optimisation problem where we minimise $|\bar{A}|$ under the constraint that $\mathcal{L}\leq\tau$ where $\tau$ is the target loss. We set the target as the loss achieved by a fully connected model. Using Lagrangian relaxation [4], we can write the min-max objective,
$\max_{\lambda>0}\min_{\theta}\quad\mathbb{E}\big{[}|\bar{A}|\big{]}+\lambda\bigg{(}\mathbb{E}\big{[}\mathcal{L}(\hat{s}^{t+1})\big{]}-\tau\bigg{)},$ (8)
where $\lambda$ is the Lagrangian multiplier. We solve this by taking gradient steps on $\theta$ and $\lambda$ alternatively. Intuitively, $\lambda$ increases the weight of the error term when the error is higher than the target. By initialising $\lambda$ to a high value, this acts as a schedule where the loss focuses on prediction accuracy when the error is high. Only once the error drops below the target are edges increasingly pruned. [^2] This is analogous to the GECO [44] framework for tuning the KL loss in a VAE. The details of this optimisation setup and other training details are included in App. A.
@@ 4 Experiments
Figure 2: Example rollouts in the two simulated environments with the learnt local causal graph visualised. Blue and red arrows between the icons indicate the identified causal edges and intervention targets respectively. In the Interventional Pong example, the environment is intervened such that the ball slows down in the middle. SPARTAN correctly identifies the same causal dependencies as the ground-truth (e.g. ball causes the paddles to follow). In contrast, the Transformer baseline learns edges that do not correspond to the ground-truth. Similarly in the CREATE example, SPARTAN learns the correct causal edges (e.g. the green ball being influenced by the plank when it bounces) while Transformer learns many spurious interventions.
@@
We evaluate our model on three domains: Interventional Pong, CREATE, and Traffic. The Interventional Pong dataset [36], a standard benchmark for causal representation learning, is based on the Pong game with interventions. The CREATE [22] environment is a 2D physics simulation which consists of interacting objects such as ladders, walls, cannons and balls. This is similar to the PHYRE dataset [2] on which SlotFormer [61] is evaluated, but has more interaction types which allows for more interventions to better showcase the capabilities of SPARTAN. For these two domains, we obtain object slot representations by providing ground-truth masks for objects and encoding each object separately using a VAE [27]. In order to evaluate our model on more realistic tasks, the Traffic domain uses the Waymo Open Dataset [52] which are collected in real life. The task is to predict the motion of the ego vehicle conditioned on the observed vehicle trajectories (i.e. past positions) and map layout lines. For evaluating the accuracy of the learnt causal graphs, we compare against ground-truth causal graphs for the two simulated domains. For the Traffic domain, we compare against human-labeled [^3] We note that these labels are subjective and only act as a qualitative reference for investigating whether the models learn meaningful edges. causal graphs [47]. In the simulated domains, we also train the models using interventional data as described in Sec. 3.2. Interventions are defined as changes to the simulated dynamics such as changing the strength of gravity. See App. B for details.
@@
In order to investigate whether SPARTAN can maintain the state-of-the-art prediction accuracy achieved by Transformer-based models, we compare against a Transformer baseline using a Transformer for learning dynamics. This can be seen as an instantiation of the SlotFormer [61] architecture with one-step context length except the slots are learnt using ground-truth object masks. For the Traffic domain, we use MTR [51], a state-of-the-art transformer-based model designed for traffic data, as the base architecture. In terms of causal discovery, [42] argue that soft attention in Transformers alone is a strong inductive bias for learning local graphs and propose a thresholding heuristic. We compare against this approach and show that the sparsity regularisation in SPARTAN is crucial in local causal discovery. To corroborate that local causal graphs are more suitable for modelling physical interactions, we further compare against the Global Graph baseline, which is based on the transition function proposed in VCD [33] and AdaRL [20] which learn a fixed global causal graph using a masked MLP. The details of the baseline models in App. A.
Table 1: Rollout prediction error and average SHD between the learned graphs and the ground-truth causal graphs. Prediction errors are based on L2 distance measured in the token space for Pong and CREATE. The average displacement error is reported for the Traffic domain. Lower is betters.
@@
4.1 Causal Discovery
The prediction accuracy of the models are shown in table 1. Overall, SPARTAN is able to make predictions at a similar level of accuracy to, and often outperform, the Transformer baseline. We investigate whether casual discovery via sparsity regularisation is required to identify correct local causal graphs in the setting of dynamics modelling. While [42] suggests that extracting causal graphs by thresholding attention patterns is sufficient in simple state-based environments, here we demonstrate that in more complex environments, learning hard attention patterns via sparsity regularisation is necessary for accurate local causal discovery. Fig. 2 visualises example rollouts with the local causal graphs. [^4] When evaluating the Transformer baseline, we pick a threshold that achieves the best structural hamming distance. Note that this is not possible in practice without access to ground-truth information. Here, we observe that SPARTAN can reliably recover the relevant causal edges and intervention targets at each timestep. In the Interventional Pong environment, our method identifies that the paddles are constantly influenced by the ball as they follow it. In contrast, the Transformer baseline gives erroneous edges such as "Score $\rightarrow$ Ball" (frame 1). Similarly in the CREATE environment, where SPARTAN accurately discovers sparse interactions between objects. In the Traffic domain, Fig. 3 shows that SPARTAN learns to attend to vehicles in neighbouring lanes, which is largely consistent with the human labels, whereas the Transformer baseline attends to many irrelevant vehicles. In this case, heuristics based on spatial proximity would also fail, since there are vehicles that are far ahead that are relevant while vehicles that are close by but are in the opposite lane, for example, should be ignored. This further highlights the importance of learning local causal graphs. Quantitatively, we evaluate the Structural Hamming Distance, a commonly used metric in graph structure learning, between the learned graphs and the ground-truth. Table 1 shows that SPARTAN achieves significantly lower distance compared to the baselines across all domains. We do not apply the Global Graph baseline to the Traffic domain since the fixed graph approach cannot be extended to include different number of objects.
@@ 4.2 Robustness and Adaptation
Figure 4: Adaptation errors on the two datasets. Each model has access to 5 trajectories with unknown environment index. "Unseen environments" refers to interventions that are not in the training set. Red lines indicates the prediction error if the environment index is provided, which serves as a lower bound. SPARTAN (blue) consistently achieves the lowest errors across environments. Remarkably, in seen environments, SPARTAN can often reach the lower bound with only 5 observed trajectories.
@@
[47] proposed to evaluate the robustness of models by considering the change of the prediction error when non-causal entities are removed from the scene. This can be seen as a special case of interventions which only targets variables that are not the cause of the predicted variable. Concretely, for each object in the scene, we remove any object that is not in the ground-truth parent set and compute the mean absolute value of the percentage change in the prediction error. Here, a model that correctly captures the causal structure would be robust whilst a model which learns spurious correlations would have large fluctuations. Table 2 shows that compared to a non-regularised Transformer, SPARTAN is significantly more robust, suggesting that the Transformer overfits to spurious correlations between entities. In the Interventional Pong domain, the ground-truth causal graph is largely consistent across time steps. As such, the Global Graph baseline is sufficient in capturing the causal structure and hence remain robust. However, this degrades in the more complex CREATE domain where the local causal graph is more state-dependent, while SPARTAN achieves the best robustness. In the Traffic domain, we also implement a local attention variant of the Transformer baseline, proposed in MTR [51], which applies a spatial K-nearest neighbour mask on the attention pattern between the tokens. Whilst this heuristic improves the model, SPARTAN remains significantly more robust to non-causal changes. A wide range of other model architectures are evaluated on the traffic dataset in [47], with reported changes ranging from 25% to 38%, which is consistent with our findings. Remarkably, the authors in [47] propose a data augmentation scheme aimed to improve model robustness resulting in a change of 22%, which is still significantly higher than what SPARTAN achieves. This further highlights the efficacy of our method.
@@
To investigate adaptation efficiency, we adapt the models on a sample of five trajectories from an intervened environment. For SPARTAN, we perform adaptation as described in Sec. 3.2. We follow a similar procedure of optimising over intervention tokens for the Global Graph baseline. [^5] This is similar to the adaptation algorithm in [20], on which the Global Graph baseline is based. For the Transformer baseline, adaptation is performed via gradient finetuning on the provided trajectories. We also adapt the models to previously unseen environments where the interventions performed are not included in the training set. In the Interventional Pong dataset, the new environments are obtained by combining interventions from the training set, e.g. intervening on the ball and the motion of the paddle at the same time. In CREATE, we change the scene composition as well as the dynamics, i.e. we change the number of objects in the scene. In this scenario, the Global Graph baseline cannot be applied. We present the results in Fig. 4, which shows that SPARTAN consistently outperform the baselines in this few-shot adaptation setting for both seen and unseen environments.
@@ 5 Related Works
Our work is motivated by recent advances in causal machine learning [48, 25]. In particular, [45] provides a theoretical account for how identifying a causal world model leads to generalisable agents. Our work can be seen as an instantiation of the Independent Causal Mechanism (ICM) principle [48], which posits that dynamics can be factorised into causal mechanisms that can be independently intervened. Inspired by ICM, RIM [16] mimics the independent causal mechanism structure by representing state dynamics as independent recurrent units. Using sparse communications between mechanisms as inductive bias, several works build on RIMs to model unitary [1] or binary interactions [34, 15, 14] between objects as a set of independent mechanisms. Our work takes an orthogonal approach of learning dynamics models using sparse regularisation. Specifically, we take inspiration from the DCDI framework [5] which formulates causal discovery as continuous optimisation. To this end, VCD [33] and AdaRL [20] which also investigate adaptability by learning a fixed global causal graph using sparsity regularisation. In contrast, our model learns state-dependent local graphs which are better suited for modelling physical interactions. The notion of local causal model is discussed in [50, 42, 21]. CAI [50] considers estimating the local influence an agent has on the environment as an exploration signal. The CoDA family of works [42, 43] demonstrates that local causal graphs enable the generation of counterfactual data for data augmentation. However, estimating a local causal model from data remains challenging. [42] proposes a thresholding heuristic for identifying causal edges from soft attention patterns. ELDEN [59] uses the partial derivative between states to infer local causal graphs. However, this requires minimising the the magnitude of partial derivatives which is computationally expensive beyond simple state-based settings. Closest to our work is FCDL [21] which also uses sparsity regularisation in the setting of learning a quantised codebook of state-dependent causal connections. In contrast to these approaches, which requires state observations, our method can operate on object-centric image embeddings. Moreover, due to the flexibility of the Transformer architecture, SPARTAN is able to learn in more realistic scenarios where the number and type of objects change between data samples.
@@
Our work is situated within the context of learning world models [e.g. 17, 18]. In particular, we build on methods that learn dynamics models over object-centric representations [12, 13, 37, 60, 11]. In this space, various architectures have been proposed to model the interactions between object slots as message passing, including GNNs [28, 53], RNN [55, 24] and pair-wise interactions [57]. SlotFormer [61] achieves state-of-the-art results by applying a Transformer-based dynamics model [56, 46, 38] on pre-trained object-slots [37]. We build on this approach by extending the Transformer architecture with learnable sparse masks and performing causal discovery on object tokens, resulting in improved interpretability and adaptability while maintaining the same level of prediction accuracy.
@@
In a broader context, enforcing sparse connections between tokens in Transformers has also been explored in other settings such as NLP [8, 7, 35] and computer vision [63]. These methods requires pre-defined mask based on domain knowledge such as sentence structures or spatial proximity, and are primarily designed for computation efficiency. SPARTAN can be seen as a way to learn these masks from data by regularising the flow of information between tokens.

### v3

@@ 3.2 Interventions
We can extend our model to represent interventions in cases where the model is trained on a dataset with a set of environments with intervened dynamics. [^1] Note that this is an extension, rather than a requirement, to our approach to capture multi-environment training. In single environment cases, SPARTAN can be directly applied as described above. During training, the model has access to the environment index $I$ for each transition but does not know which object is affected by the intervention. Conditioned on the current state $s^{t}_{1:N}$ and the environment index $I$, the model needs to identify a subset of $s^{t+1}_{1:N}$ as intervention targets and change the predictions (cf. Eq. 2). To do this, we keep a learnable codebook of intervention tokens, $\mathcal{T}_{1:K}\in\mathbb{R}^{d\times k}$, where $d$ is the dimension of the state tokens and $K$ is the number of interventions. We append the corresponding intervention token $\mathcal{T}_{I}$ to the object tokens and proceed as described in the previous subsection. Here, $\bar{A}_{i(N+1)}=0$ indicates that there is no path from the intervention token to $s^{t+1}_{i}$ and therefore $s^{t+1}_{i}$ is not an intervention target, i.e., the prediction for $s^{t+1}_{i}$ is not affected by the changes in this environment.
@@
Figure 2: Example rollouts in the two simulated environments with the learnt local causal graph visualised. Blue and red arrows between icons indicate the learnt causal edges and intervention targets respectively. In the Interventional Pong example, the intervention is that the ball slows down in the middle. SPARTAN correctly identifies the same causal dependencies as the ground-truth (e.g. ball causes the paddles to follow). The Transformer baseline learns edges that do not correspond to the ground-truth. Similarly in CREATE, SPARTAN learns the correct causal edges (e.g. green ball bounces off the blue plank) while Transformer learns many spurious interventions.
@@ 3.3 Training
Our aim is to fit the data distribution with the sparsest possible model in terms of causal edges and interventions. The training objective is to minimise the expected loss with sparsity regularisation,
@@
where $\theta$ are the model parameters, including the parameters for the Transformer and the intervention tokens, $\mathcal{L}$ is a loss function such as MSE, and $\lambda_{s}$ denotes the regularisation hyperparameter. Note that due to the intervention token being part of the input during training, $|\bar{A}|$ is the sum of the number of causal edges and the intervention targets. In practice, the model is sensitive to the choice of $\lambda$ since a high $\lambda$ can lead to mode collapse. We alleviate this via Lagrangian relaxation, which schedules the regularisation weight. The details of this setup and other training information are in App. A.
@@ 4 Experiments
We evaluate our model in three domains: Interventional Pong, CREATE, and Traffic. The Interventional Pong dataset [lippe2022citris], a standard benchmark for causal representation learning, is based on the Pong game with interventions. The CREATE [jain2020general] environment is a 2D physics simulation that consists of interacting objects such as ladders, cannons, and balls. This is similar to the PHYRE dataset [bakhtin2019phyre] on which SlotFormer [wu2023slotformer] is evaluated, but it has more interaction types, allowing for more interventions to better showcase the capabilities of SPARTAN. For these two domains, we obtain object slot representations by providing ground-truth masks for objects and encoding each object separately using a VAE [kingma2022autoencoding]. In order to evaluate our model on more realistic tasks, the Traffic domain uses the Waymo Open Dataset [waymo] which is collected in real life. The task is to predict the motion of the ego vehicle given the observed vehicle trajectories (i.e., past positions) and map layout lines. For evaluating the accuracy of the learnt causal graphs, we compare them against ground-truth causal graphs for the two simulated domains. For the Traffic domain, we compare against human-labelled causal graphs [causal_agents]. In the simulated domains, we also train the models using interventional data, as described in Sec. 3.2. Interventions are defined as changes to the simulated dynamics such as changing the strength of gravity. See App. B for details.
@@
In order to investigate whether SPARTAN can maintain the state-of-the-art prediction accuracy achieved by Transformer-based models, we compare our method against a Transformer baseline using a Transformer for learning dynamics. This can be seen as an instantiation of the SlotFormer [wu2023slotformer] architecture with a one-step context length, except that the slots are learnt using ground-truth object masks. For the Traffic domain, we use MTR [shi2022motion], a state-of-the-art transformer-based model designed for traffic data as the base architecture. In terms of causal discovery, existing methods [ELDEN, hwang2024finegrained] only operate on state observations where each state is represented as a scalar and are therefore not applicable in our setup, which observes object embeddings. We provide a more detailed discussion on the applicability of these prior approaches and offer extra experimental comparisons in App. C. Closest to our work is CoDA [pitis2020count], which argues that soft attention in Transformers alone is sufficient for learning local graphs and proposes a thresholding heuristic. We compare against this approach and show that sparsity regularisation in SPARTAN is crucial in local causal discovery. To corroborate that local causal graphs are more suitable for modelling physical interactions, we further compare against the Global Graph baseline, which is based on the transition function proposed in VCD [lei2023variational] and AdaRL [huang2021adarl], which learns a fixed global causal graph using masked MLPs. The details of the baseline models are provided in App. A.
Table 1: Rollout prediction error and average SHD between the learned graphs and the ground-truth causal graphs. Lower is better.
@@
Figure 4: Adaptation errors on the two datasets. Each model has access to 5 trajectories with unknown environment index. "Unseen environments" refers to interventions that are not in the training set. Red lines indicates the prediction error if the environment index is provided, which serves as a lower bound. SPARTAN (blue) consistently achieves the lowest errors across environments.
4.1 Graph Learning
The prediction accuracy of the models is shown in table 1. Overall, SPARTAN is able to make predictions at a similar level of accuracy to and often outperforms the Transformer baseline. We investigate whether casual discovery via sparsity regularisation is required to identify correct local causal graphs in the setting of dynamics modelling. While [pitis2020count] suggests that extracting causal graphs by thresholding attention patterns is sufficient in simple environments, here we demonstrate that in more complex environments, learning hard attention patterns via sparsity regularisation is necessary for accurate local causal discovery. Fig. 2 visualises example rollouts with the local causal graphs. [^2] When evaluating the Transformer baseline, we pick a threshold that achieves the best structural hamming distance. Note that this is not possible in practice without access to ground-truth information. Here, we observe that SPARTAN can reliably recover the relevant causal edges and intervention targets at each timestep. In the Interventional Pong environment, our method identifies that the paddles are constantly influenced by the ball as they follow it. In contrast, the Transformer baseline gives erroneous edges such as "Score $\rightarrow$ Ball" (frame 1). Similarly, in the CREATE environment, SPARTAN accurately discovers sparse interactions between objects. In the Traffic domain, Fig. 3 shows that SPARTAN learns to attend to vehicles in neighbouring lanes, which is largely consistent with the human labels, whereas the Transformer baseline attends to many irrelevant vehicles. In this case, heuristics based on spatial proximity would also fail since there are vehicles that are far ahead that are relevant while vehicles that are close by but in the opposite lane, for example, should be ignored. This further highlights the importance of learning local causal graphs. Quantitatively, we evaluate the Structural Hamming Distance, a commonly used metric in graph structure learning, between the learned graphs and the ground-truth. Table 1 shows that SPARTAN achieves significantly lower distance compared to the baselines across all domains. We do not apply the Global Graph baseline to the Traffic domain since the fixed graph approach cannot be extended to include a different number of objects.
@@ 4.2 Robustness and Adaptation
[causal_agents] proposed evaluating the robustness of models by considering the change in the prediction error when non-causal entities are removed from the scene. This can be seen as a special case of interventions that target only variables that are not the causes of the predicted variable. Concretely, for each object in the scene, we remove any object that is not in the ground-truth parent set and compute the mean absolute value of the percentage change in the prediction error. Here, a model that correctly captures the causal structure would be robust, whereas a model learning spurious correlations would have large fluctuations. Table 2 shows that, compared to a non-regularised Transformer, SPARTAN is significantly more robust, suggesting that the Transformer overfits to spurious correlations between entities. In the Interventional Pong domain, the ground-truth causal graph is largely consistent across time steps. As such, the Global Graph baseline is sufficient in capturing the causal structure and hence remains robust. However, this degrades in the more complex CREATE domain, where the local causal graph is more state-dependent, while SPARTAN achieves the best robustness. In the Traffic domain, we also implement a local attention variant of the Transformer baseline, proposed in MTR [shi2022motion], which applies a spatial K-nearest neighbour mask on the attention pattern between the tokens. Whilst this heuristic improves the model, SPARTAN remains significantly more robust to non-causal changes. A wide range of model architectures is evaluated in [causal_agents], with reported changes ranging from 25% to 38%, which is consistent with our findings. Remarkably, the authors in [causal_agents] propose a data augmentation scheme using ground-truth information, resulting in a change of 22%, which is still significantly worse than what SPARTAN achieves, highlighting the efficacy of our method.
@@
To investigate adaptation efficiency, we train the models to adapt to a sample of five trajectories from an intervened environment. For SPARTAN, we perform adaptation as described in Sec. 3.2. We follow a similar procedure of optimising over intervention tokens for the Global Graph baseline. For the Transformer baseline, adaptation is performed via gradient finetuning on the provided trajectories. We also adapt the models to previously unseen environments where the interventions performed are not included in the training set. In the Interventional Pong dataset, the new environments are obtained by combining interventions from the training set, e.g. intervening on the ball and the motion of the paddle at the same time. In CREATE, we change the scene composition as well as the dynamics, i.e. we change the number of objects in the scene. In this scenario, the Global Graph baseline cannot be applied. We present the results in Fig. 4, which shows that SPARTAN consistently outperforms the baselines in this few-shot adaptation setting.
@@ 5 Related Works
Our work is situated within the context of learning world models [ha2018recurrent, hafner2020dreamerv2]. In particular, we build on methods that learn dynamics models over object-centric representations [engelcke2021genesisv2, slot_attention, wu2023obpose]. In this space, various architectures have been proposed to model the interactions between object slots as message passing, including GNNs [kipf2019contrastive, tacchettiRelationalForwardModels2018], RNN [van2018relational, JiangJanghorbaniDeMeloAhn2020SCALOR] and pair-wise interactions [veerapaneni2020entity]. SlotFormer [wu2023slotformer] achieves state-of-the-art results by applying a Transformer-based dynamics model on pre-trained object-slots [slot_attention]. We build on this approach by extending the Transformer architecture with learnable sparse masks, resulting in improved robustness, adaptability, and interpretability without compromising prediction performance or the flexibility to model varying numbers of objects.
Our approach is motivated by recent advances in causal machine learning [scholkopf2021toward, kaddour2022causal]. In particular, [richens2024robust] provides a theoretical account of how identifying a causal world model leads to generalisable agents. Concretely, our sparsity regularisation-based approach is motivated by optimisation-based causal discovery methods [brouillard2020differentiable] that use sparsity as an inductive bias for learning causal relationships. In this vein, VCD [lei2023variational] and AdaRL [huang2021adarl] learn fixed causal graphs using sparsity regularisation in the context of world models. We extend these approaches by learning state-dependent local graphs that are better suited for modelling physical interactions. The notion of local causal model for dynamics is discussed in [pitis2022mocoda, pitis2020count, ELDEN], which demonstrates the potential benefits of local causal models in terms of data-efficiency, robustness, and exploration. The CoDA family of works [pitis2020count, pitis2022mocoda, ameperosa2024rocoda] uses local causal graphs to enable the generation of counterfactual data for data augmentation. However, estimating a local causal model from data remains challenging. Several approaches to learning local causal structures have been proposed. CAI [seitzer2021causal] uses conditional mutual information to estimate the local influence that an agent has on the environment. ELDEN [ELDEN] uses sparsity-regularised partial derivatives between states based on a learnt dynamics model to infer local causal graphs. FCDL [hwang2024finegrained] uses sparsity regularisation to learn a quantised codebook of possible causal connections. The key difference between these approaches and ours is that they require state observations, whereas our method can operate on object-centric image embeddings. Moreover, owing to the flexibility of transformers, SPARTAN is able to learn in more realistic scenarios where the number of objects or the scene composition changes between data samples. We provide a more detailed discussion on the applicability of existing methods in our setting and present additional experiments in App. C. Closest to our work is [pitis2020count] which uses attention patterns in transformer layers as causal edges. We empirically show that our model achieves significantly improved accuracy in terms of the discovered graph structure.
@@
In a broader context, enforcing sparse connections between tokens in Transformers has also been explored in other settings, such as NLP [correia2019adaptively, child2019generating, leviathan2024selectiveattentionimprovestransformer] and computer vision [zhou2024adapt]. These methods require pre-defined masks based on domain knowledge, such as paragraph structures. SPARTAN differs from these approaches as it does not require any pre-defined masks and can be seen as a way to learn these masks from data. To this end, investigating the application of sparsity-regularised hard attention in these domains outside of world modelling offers a tantalising avenue for future research.
On a conceptual level, our work is also related to the notion of the Consciousness Prior [bengio2017consciousness], which advocates for an attention-like mechanism that selects a sparse bottleneck of ’active’ entities in any given scene. This idea is realised in [zhao2021a], showing that the sparsity induced by the Consciousness Prior leads to improved generalisation in model-based planning tasks. Our approach differs from this idea in two principal ways: 1) [zhao2021a] requires pre-specifying the number of active entities, whereas the sparse attention in our approach flexibly attends to a varying number of objects in a state-dependent manner. 2) In the Consciousness Prior framework, the model is sparse in the sense that it chooses a small subset of objects that is active. In contrast, we consider a complementary kind of sparsity, where we model the dynamics of all objects but use sparse attention to determine how each object depends on other objects.